# ScaffoldGPT: A Scaffold-based Large Language Model for Drug Improvement

### v1

@@ 5.1 Experimental configuration
ScaffoldGPT adopts APO finetuning as the second step, following the completion of pretraining the LLM-based generator. Table 1 demonstrates the effectiveness of APO through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o Top-N), which shows that after applying APO finetuning, performance improved on most properties. Additionally, ScaffoldGPT vs. ScaffoldGPT (w/o APO) validates the importance of the APO component. By applying APO on top of pretraining and Top-N decoding, performance improved. Both cases demonstrate the effectiveness of APO finetuning.
@@
ScaffoldGPT adopts the Top-N decoding strategy as the final step followed by APO finetuning. Table 1 demonstrates the effectiveness of Top-N through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o APO), showing that after applying the Top-N decoding strategy on top of pretrained LLM, performance improved across most properties. Moreover, ScaffoldGPT vs. ScaffoldGPT (w/o Top-N) illustrates that after applying APO on top of pretraining and RL, performance still improves on multiple attributes, surpassing all baselines. Furthermore, by comparing ScaffoldGPT (w/o APO) and ScaffoldGPT (w/o Top-N), we observe that applying Top-N decoding alone enhances performance more than applying APO alone.
@@
When removing Top-N component, we employ multinomial generation, where multinomial sampling randomly selects the next token from the entire vocabulary based on the model’s probability distribution. In the ablation study detailed in Table 4, we examined Top-N over Top-P, Top-K, and Top-PK with K=20 and P=0.95. The results indicate that Top-N surpasses the other strategies in most metrics.
@@ 6 Conclusion
In this study, we introduce ScaffoldGPT, a novel framework named designed for drug optimization. It incorporates a unique scaffold-based LLM design, a three-step optimization process, a two-phase incremental training method, and a novel Top-N decoding strategy, facilitating controlled reward-guided generation using pretrained LLMs. To showcase the superior performance of ScaffoldGPT, we conduct evaluations on real-world viral and cancer-related datasets, comparing it against eight competing baselines. Our results demonstrate that ScaffoldGPT surpasses all baselines across the majority of performance metrics, underscoring its efficacy. Our work highlights ScaffoldGPT’s effectiveness in drug optimization, as evidenced by enhancements in various pharmaceutical properties. Currently, ScaffoldGPT is limited to handling molecules in SMILES format. We are looking into expanding ScaffoldGPT’s capabilities to accommodate a broader range of drug representation formats as our future direction.
Acknowledgements
This work is supported in part by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.
Impact Statement
This paper explores advancements in machine learning for developing language models aimed at drug optimization. While language models for drug optimization may have certain societal implications, we do not find any that must be specific emphasis here.
@@ A.1 Pre-training and finetuning dataset
For pretrainig, We used the ZINC dataset, filtering for Standard, In-Stock, and Drug-Like molecules, resulting in approximately 11 million molecules. For preprocessing, we perform a few straightforward steps: 1. We first canonicalize smiles Chem.MolToSmiles(Chem.MolFromSmiles(mol),True). 2. We filter out molecules whose scaffold SMILES is an empty string. These preprocessing steps are also included in the huggingface data repo.
For finetuning, we utilize 1 million compounds from the ZINC15 dataset, docked to the 3CLPro protein (PDB ID: 7BQY) linked to SARS-CoV-2 and the RTCB protein (PDB ID: 4DWQ) associated with human cancer, as sourced from the latest Cancer and COVID dataset by Liu et al. [22], across all baselines.
@@ A.3 BPE Tokenization
By iteratively merging the most frequent pairs, BPE builds a robust vocabulary that captures common subword units, allowing for more efficient and flexible text representation.
@@ A.4 Surrogate model
The surrogate model [44] is a simplified variant of a BERT-like transformer, extensively utilized in natural language processing. In this model, tokenized SMILES strings are inputted and then embedded with positional information. The resulting outputs are subsequently fed into a series of five transformer blocks, each comprising a multi-head attention layer (21 heads), a dropout layer, layer normalization with residual connection, and a feedforward network. This feedforward network consists of two dense layers followed by dropout and layer normalization with residual connection. Following the stack of transformer blocks, a final feedforward network is employed to generate the predicted docking score. The validation $r^{2}$ values are 0.842 for 3CLPro and 0.73 for the RTCB dataset.
@@ A.5 Drug Optimization illustration on COVID benchmark
A.6 Binding sites of 3clpro and RTCB
(a) 3CLPro.
(b) RTCB.
Figure 2: The binding sites of proteins 3CLPro (PDB ID: 7BQY) (Left) and RTCB (PDB ID: 4DWQ) (Right). Binding sites are defined around the crystallized compound using Open Eye software.
A.7 Computing infrastructure and wall-time comparison
We trained our docking surrogate models using 4 nodes of the supercomputer where each node contains CPUs (64 cores) and 4 A100 GPU nodes [8]. The training time for each model was approximately 3 hours.
@@ A.7 Computing infrastructure and wall-time comparison
A.8 Hyperparameters and architectures
Table 7 and 8 provides a list of hyperparameter settings we used for our experiments.
For APO finetuning and experimentation, 1280 molecules were selected from each of the RTCB and 3CLPro datasets, with docking scores ranging from -14 to -6. This range is based on [23].

### v2

@@ 5.1 Experimental configuration
ScaffoldGPT adopts APO finetuning as the second step, following the completion of pretraining the GPT-based generator. Table 1 demonstrates the effectiveness of APO through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o Top-N), which shows that after applying APO finetuning, performance improved on most properties. Additionally, ScaffoldGPT vs. ScaffoldGPT (w/o APO) validates the importance of the APO component. By applying APO on top of pretraining and Top-N decoding, performance improved. Both cases demonstrate the effectiveness of APO finetuning. Note that APO may compromise certain reward metrics, such as similarity or validity, if this trade-off leads to improved performance in the target weighted objective.
@@
ScaffoldGPT adopts the Top-N decoding strategy as the final step followed by APO finetuning. Table 1 demonstrates the effectiveness of Top-N through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o APO), showing that after applying the Top-N decoding strategy on top of pretrained GPT, performance improved across most properties. Moreover, ScaffoldGPT vs. ScaffoldGPT (w/o Top-N) illustrates that after applying APO on top of pretraining and RL, performance still improves on multiple attributes, surpassing all baselines. Furthermore, by comparing ScaffoldGPT (w/o APO) and ScaffoldGPT (w/o Top-N), we observe that applying Top-N decoding alone enhances performance more than applying APO alone.
@@
Following the setup described in §5.1, we perform an ablation study on sampling strategies. When removing Top-N component, we employ multinomial generation, where multinomial sampling randomly selects the next token from the entire vocabulary based on the model’s probability distribution. In the ablation study detailed in Table 2, we examined Top-N over Top-P, Top-K, and Top-PK with K=20 and P=0.95. The results indicate that Top-N surpasses the other strategies in most metrics.
Original             Scaffold REINVENT ScaffoldGPT 1 ScaffoldGPT 2 ScaffoldGPT 3
Molecule [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE String Cc1ccc(O)c(Nc2nnc (-c3ccccc3)c(=O)[nH]2)c1 CCOc1ccc(O)c(Nc2nnc (-c3ccccc3)c(=O)[nH]2)c1 Cc1ccc(-c2nnc(Nc3ccc (C(C)C)cc3)[nH]c2=O)cc1 COc1ccc(-c2nnc(Nc3cccc (C)c3)[nH]c2=O)cc1 Cc1ccc(-c2nnc(Nc3cccc (C(C)C)c3)[nH]c2=O)cc1
Scaffold O=c1[nH]c(Nc2cc ccc2)nnc1-c1ccccc1 c1ccc(Nc2nnc(-c3 ccccc3)[nH]2)cc1 same as original same as original same as original
Similarity - 0.878 0.826 0.911 0.866
Docking ()($\downarrow$)( ↓ ) -10.031 -9.258 -11.478 $\checkmark$ -11.474 $\checkmark$ -11.087 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.646 0.624 0.762 $\checkmark$ 0.774 $\checkmark$ 0.762 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 2.390 2.417 2.298 $\checkmark$ 2.257 $\checkmark$ 2.356 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 2.590 2.680 $\checkmark$ 4.007 $\checkmark$ 2.893 $\checkmark$ 4.007 $\checkmark$
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.618 0.589 0.759 $\checkmark$ 0.753 $\checkmark$ 0.754 $\checkmark$
Table 3: One optimization example from cancer benchmark. Every generated molecules retains the scaffold, with all desired properties improved compared to the original. represents the top-priority target objective. $\checkmark$ indicates improved property.
@@ 6 Conclusion
We have introduced ScaffoldGPT, a novel framework for drug optimization. This framework incorporates a unique scaffold-based GPT design, a three-step optimization process, a two-phase incremental training method, and a novel Top-N decoding strategy that facilitates controlled reward-guided generation using GPT. To showcase the superior performance of ScaffoldGPT, we conduct evaluations on real-world viral and cancer-related datasets to compare it against eight competing baselines. Our results demonstrate that ScaffoldGPT surpasses all baselines across the majority of performance metrics, underscoring its efficacy. Our work highlights ScaffoldGPT’s effectiveness in drug optimization, as evidenced by enhancements in various pharmaceutical properties. Currently, ScaffoldGPT is limited to handling molecules in SMILES format. We are working to expand ScaffoldGPT’s capabilities to accommodate a broader range of drug representation formats as a future direction.
@@ A.1 Pre-training and finetuning dataset
For pretrainig, We used the ZINC dataset, filtering for Standard, In-Stock, and Drug-Like molecules, resulting in approximately 11 million molecules. For preprocessing, we perform a few straightforward steps:
1. We first canonicalize smiles Chem.MolToSmiles(Chem.MolFromSmiles(mol),True).
2. We filter out molecules whose scaffold SMILES is an empty string. These preprocessing steps are also included in the huggingface data repo.
For finetuning, we utilize 1 million compounds from the ZINC15 dataset, docked to the 3CLPro protein (PDB ID: 7BQY) linked to SARS-CoV-2 and the RTCB protein (PDB ID: 4DWQ) associated with human cancer, as sourced from the latest Cancer and COVID dataset by Liu et al. (2023a), across all baselines.
@@ A.3 BPE Tokenization
BPE constructs a robust vocabulary by iteratively merging the most frequent token pairs, effectively capturing common subword units for more efficient and flexible text representation. The resulting vocabulary comprises 3,152 tokens and includes special tokens as well. For instance, the sequence <L> is tokenized into three separate tokens: $<$, L, and $>$. The tokenizer was trained on 10 million molecules from the ZINC dataset, ensuring comprehensive coverage of chemical elements.
@@ A.4 Surrogate model
The surrogate model (Vasan et al., 2023) is a simplified variant of a BERT-like transformer, extensively utilized in natural language processing. In this model, tokenized SMILES strings are inputted and then embedded with positional information. The resulting outputs are subsequently fed into a series of five transformer blocks, each comprising a multi-head attention layer (21 heads), a dropout layer, layer normalization with residual connection, and a feedforward network. This feedforward network consists of two dense layers followed by dropout and layer normalization with residual connection. Following the stack of transformer blocks, a final feedforward network is employed to generate the predicted docking score. The validation $r^{2}$ values are 0.842 for 3CLPro and 0.73 for the RTCB dataset.
@@ A.5 Drug Optimization illustration on COVID benchmark
A.6 Computing infrastructure and wall-time comparison
We trained our docking surrogate models using 4 nodes of the supercomputer where each node contains CPUs (64 cores) and 4 A100 GPU nodes (Facility, ). The training time for each model was approximately 3 hours.
@@ A.6 Computing infrastructure and wall-time comparison
A.7 Hyperparameters and architectures
Table 6 and 7 provides a list of hyperparameter settings we used for our experiments.
For APO finetuning and experimentation, 1280 molecules were selected from each of the RTCB and 3CLPro datasets, with docking scores ranging from -14 to -6. This range is based on (Liu et al., 2024b).